OpenStudy (anonymous):

Two times a number is divided by 3 more than the number. If the result is 7, then what was the original number? Two times a number is divided by 3 more than the number. If the result is 7, then what was the original number? @Mathematics

OpenStudy (anonymous):

2x/(3+x)=7

OpenStudy (anonymous):

right? no?

OpenStudy (anonymous):

whats the answer then?

OpenStudy (hoblos):

2x/x+3 =7 2x = 7x +21 5x =-21 => x=- 21/5